OpenStudy (anonymous):

What value of y makes the equation true? (–9.7 • 24) • (–0.25) = –9.7 • (24 • y)

OpenStudy (anonymous):

I'll give a medal :)

OpenStudy (tkhunny):

Have you considered evaluating the various expressions and solving for y?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

.................

OpenStudy (anonymous):

CAN SOMEONE HELP PLEASE :)

OpenStudy (tkhunny):

Demonstrate it. (–9.7 • 24) • (–0.25) = –9.7 • (24 • y) Divide by -9.7 (24) • (–0.25) = (24 • y) Are you SURE you want us to believe that you couldn't do that? Okay, what else?

OpenStudy (anonymous):

uuhhh not really understanding with those question marks

OpenStudy (tkhunny):

Refresh the screen. They usually go away.

OpenStudy (gorv):

it is similr to (x*y)*z=x*(y*z) just a shift of bracket

OpenStudy (gorv):

it is a property

OpenStudy (gorv):

ascociative property

OpenStudy (tkhunny):

Let's try this... (–9.7*24)*(–0.25) = –9.7*(24*y) Divide by -9.7 (24)*(–0.25) = (24*y) Are you SURE you want us to believe that you couldn't do that? Okay, what else?

OpenStudy (anonymous):

whats the anwser jk

OpenStudy (gorv):

@tkhunny itzzz associative property man

OpenStudy (anonymous):

no they asked me whats the value of y

OpenStudy (anonymous):

Thx :)

OpenStudy (tkhunny):

Yes, you can use the associative property, but CuteGirl903k doesn't seem to like that, either. Feel free to suggest a 3rd approach. Did you divide by -9.7? That would help you on your way. Do you know the associative property of multiplication? That will make this an eyeball problem. Just look at it and read off the answer.